A battery is connected across a series combination of two identical resistors. If the potential difference across the terminals is V and the current in the battery is i, then:

Which of these statement are true?

The potential difference across each resistor is V and the current in each resistor is i

The potential difference across each resistor is V/2 and the current in each resistor is i

The potential difference across each resistor is V and the current in each resistor is i/2

The potential difference across each resistor is V/2 and the current in each resistor is i/2

None of these are true
Answer #1

Cureent passes through the resisters is same (i).Because the two resisters are in series.

Resistance of each resistor = r

Effective resistance R = r + r = 2r

Potential difference V = iR

Potential difference through each resistor V' = ir = i(R/2) = (iR)/2  = V/2

So The potential difference across each resistor is V/2 and the current in each resistor is i